# TOB-SVD: Total-Order Broadcast with Single-Vote Decisions in the Sleepy Model

### v2

@@ 4.1 Starting point
The protocol still relies on the key ideas of the time-shifted quorum technique presented above. A notable difference is that grade 00 outputs in MR [10] are computed at time $3\Delta$ using $\mathsf{VOTE}$ messages, whereas our GA protocol does not have any other message other than $\mathsf{LOG}$ messages, and it computes grade 00 outputs at time $2\Delta$. That said, votes in their GA are cast (almost) in the same way as grade 00 outputs are computed in our protocol, and they are essentially used to propagate the information forward until the end of the protocol. Another difference is the use of $|V_{\Lambda}^{\Delta}\cap V_{\Lambda}^{{3\Delta}}|$ when determining outputs of grade $1$, which is related to the treatment of equivocations. Let us initially ignore that, and pretend that we output $(\Lambda,1)$ when $|V_{\Lambda}^{\Delta}|>|S^{3\Delta}|/2$.
@@
To ensure that the supporting $\mathsf{LOG}$ messages used when attempting to output a log $\Lambda$ with grade 00 are more than when doing so with grade $1$, we use $V^{\Delta}_{\Lambda}\cap V^{3\Delta}_{\Lambda}$ instead of just $V^{\Delta}_{\Lambda}$. Crucially, any validator $v_{k}$ which is seen as an equivocator by validator $v_{j}$ at time $2\Delta$ will also be seen as an equivocator by validator $v_{i}$ at time $3\Delta$, since $v_{j}$ would forward the equivocating $\mathsf{LOG}$ messages of $v_{k}$ upon receiving them. In particular, logs sent by $v_{k}$ will not be contained in $V^{3\Delta,i}$, and thus also not in $V^{\Delta,i}_{\Lambda}\cap V^{3\Delta,i}_{\Lambda}$. This ensures that $V^{\Delta,i}_{\Lambda}\cap V^{3\Delta,i}_{\Lambda}\subseteq V^{2\Delta,j}_{\Lambda}$, since any log in $V^{\Delta,i}\setminus V^{2\Delta,j}$ is an equivocation, which implies that it is absent from $V^{3\Delta,i}$. In other words, we separate the initial determination of possible supporting $\mathsf{LOG}$ messages from the final determination of which $\mathsf{LOG}$ messages should be treated as equivocations and removed, and we apply the time-shifted quorum technique to the set of equivocators as well: like the perceived participation level, the set of equivocators increases when going from the output phase for grade 00 to the output phase for grade $1$.
@@
Our Graded Agreement with grades $k=3$. Building upon our Graded Agreement protocol with $k=2$ grades, we extend it to a Graded Agreement with $k=3$ grades by applying the time-shifted quorum technique twice. The first application happens during time $[2\Delta,4\Delta]$, and ensures the Graded Delivery property between grades 00 and $1$, exactly in the same way as in our Graded Agreement with $k=2$ grades. This application is nested inside the second one, which ensures the Graded Delivery property between grades $1$ and $2$, and happens during time $[\Delta,5\Delta]$. Overall, the relevant inclusions are $V^{\Delta}_{\Lambda}\cap V^{5\Delta}_{\Lambda}\subseteq V^{2\Delta}_{\Lambda}\cap V^{4\Delta}_{\Lambda}\subseteq V^{3\Delta}_{\Lambda}$, for any $\Lambda$, and $S^{3\Delta}\subseteq S^{4\Delta}\subseteq S^{5\Delta}$ (each set can belong to a different validator). The protocol is given in Figure 2.
@@
Upon input $\Lambda$, a validator $v_{i}$ runs the following algorithm whenever awake. All validators awake at time $3\Delta$ participate in the output phase for grade 00. A validator participates in the output phase for grade $1$ at time $4\Delta$ if and only if it was awake also at time $2\Delta$. A validator participates in the output phase for grade $2$ at time $5\Delta$ if and only if it was awake also at time $\Delta$. At any time, honest validators forward any message received. Up to two different $\mathsf{LOG}$ messages per sender are forwarded upon reception. 1. Input phase, $(t=0)$: Broadcast ${\langle\mathsf{LOG},\Lambda\rangle}_{i}$. 2. $(t=\Delta)$: Store $V^{\Delta}$. 3. $(t=2\Delta)$: Store $V^{2\Delta}$. 4. Output phase for grade 00, $(t=3\Delta)$: If $|V^{3\Delta}_{\Lambda}|>|S^{3\Delta}|/2$, output $(\Lambda,0)$. 5. Output phase for grade $1$, $(t=4\Delta)$: If awake at time $2\Delta$: If $|V^{2\Delta}_{\Lambda}\cap V^{4\Delta}_{\Lambda}|>|S^{4\Delta}|/2$, output $(\Lambda,1)$. 6. Output phase for grade $2$, $(t=5\Delta)$: If awake at time $\Delta$: If $|V^{\Delta}_{\Lambda}\cap V^{5\Delta}_{\Lambda}|>|S^{5\Delta}|/2$, output $(\Lambda,2)$.
@@
Our main result: TOB-SVD Byzantine Total-Order Broadcast protocols can be built using Graded Agreement primitives, with some protocols invoking Graded Agreement multiple times for a single decision [10, 8, 9]. At its core, dynamically available TOB aims to guarantee the total order delivery of messages among honest validators, even in the face of challenges like dynamic participation. However, existing TOB protocols [10, 8, 9, 7] have grappled with challenges related to latency [10], resilience [8], and scalability [7]. Multiple invocations of GA for a single decision, meaning multiple voting phases per new block, might exacerbate scalability and latency issues, especially when considering implementation in blockchain networks with hundreds of thousands of participants weighing in on every decision.
@@
Specifically, our TOB protocol (Section 6), which works in the $(5\Delta,2\Delta,1)$-sleepy model, proceeds over a series of views, each spanning a duration of $4\Delta$. Every view $v$ initiates a Graded Agreement $GA_{v}$ with grades 00, $1$, and $2$, which extends and overlaps with the following $GA_{v+1}$ during view $v+1$. This structure implies that a single view does not encapsulate a full cycle of a Graded Agreement; instead, a $GA_{v}$ initiated in view $v$ concludes its operations only in the succeeding view $v+1$.
@@
Figure 3: In the middle, views $v-1$, $v,$ and $v+1$ of our Total-Order Broadcast protocol, each with its three phases. At the top and bottom, respectively, $GA_{v-1}$ and $GA_{v}$ – the Graded Agreement protocols that are run as part of the TOB protocol. Arrows indicate that outputs of a GA are used by a parallel phase of the TOB and/or the next GA: outputs of grade 00 of $GA_{v-1}$ are extended by proposals of view $v$, outputs of grade $1$ of $GA_{v-1}$ are extended by votes in the TOB, which exactly correspond to the inputs to $GA_{v}$, while outputs of grade $2$ of $GA_{v-1}$ are decided in view $v$.
At a high level, our protocol operates through the following phases, depicted in Figure 3.
  1. 1.
    Proposal Phase: This phase corresponds to the grade 00 output phase of $GA_{v-1}$. Validators propose a log extending their candidate log, which is the grade 00 output from $GA_{v-1}$.
  2. 2.
    Voting Phase: This phase corresponds to both the input phase of $GA_{v}$ and the grade $1$ output phase of $GA_{v-1}$. Validators with a grade $1$ output treat it as a lock. To maintain Safety, they only input to $GA_{v}$ either the lock or a proposal that extends it.
  3. 3.
    Decision Phase: This phase corresponds to the grade $2$ output phase of $GA_{v-1}$. Validators decide their grade $2$ output.
In an execution of our Total-Order Broadcast protocol (Figure 3), validators start the $GA$ of view $v$ ($GA_{v}$) at time $t=\Delta$ by inputting the outputs of grade 1 from $GA_{v-1}$. For a validator to successfully output a log with grade 1 from $GA_{v-1}$, it is required that this validator was awake at $t=2\Delta$ within $GA_{v-1}$, which aligns with $t=3\Delta$ in view $v-1$ of the Total-Order Broadcast. For this reason, we need to assume a stabilization period of $2\Delta$, where participation remains stable, in order to ensure that enough validators successfully start $GA_{v}$. We expand on this in Section 6.
@@
On the number of grades. Our protocol requires $k=3$ grades to reach a decision with a single instance of the GA. Specifically, grade 2 enables an instant decision, which would otherwise necessitate two instances of GA, each with $k=2$ grades. For comparison, in constructions like 1/3MMR, participants first input their logs into the first GA. Afterward, they input the longest log output with grade 1 from the first GA into the second GA. The decision is then made from the grade 1 output of the second GA, which corresponds to the grade 2 output in our GA with $k=3$ grades.
@@
It is important to note that the decision phase of our TOB protocol is solely focused on making decisions and nothing else. By omitting the decision phase, the protocol can be simplified, still maintaining one GA per decision. However, this modification would shift the protocol from ensuring deterministic Safety to probabilistic Safety. As a result, it would align more closely with a different line of research focused on probabilistically safe, dynamically available protocols, such as those presented in [4, 5]. Nevertheless, exploring probabilistically safe TOB protocols falls outside the scope of this work.
5 Analysis of Graded Agreement protocols
In this section, we analyse the two Graded Agreement protocols presented in Section 4.
@@ 5.1 Graded Agreement with $k=2$ grades
We begin by analysing the Graded Agreement protocol with $k=2$ grades.
The protocol lasts $3\Delta$ time, and requires $|H_{0,0,3\Delta}|>|B_{3\Delta}|$, i.e., that the validators which are initially awake and honest throughout the execution outnumber all adversarial validators. Therefore, it specifically works in the $(3\Delta,0,1)$-sleepy model. Honest validators which are awake during the output phase for grade 1, at time $t=3\Delta$, only participate in it if they were also awake at time $t=\Delta$.
@@
Let $m=(\Lambda^{\prime\prime},v_{k})$ for some $\Lambda^{\prime\prime}\succeq\Lambda$ and some validator $v_{k}$, and say $m\in V^{\Delta,i}\cap V^{3\Delta,i}$. Since $m\in V^{\Delta,i}$, validator $v_{i}$ forwards message ${\langle\mathsf{LOG},\Lambda^{\prime\prime}\rangle}_{k}$ at time $\Delta$ and $v_{j}$ receives it by time $2\Delta$. Therefore, either $m\in V^{3\Delta,j}$ or $v_{k}\notin V^{3\Delta,j}$, because receiving any log other than $\Lambda^{\prime\prime}$ from $v_{k}$ would lead to $v_{j}$ having equivocation evidence for $v_{k}$. Since $m$ is for $\Lambda^{\prime\prime}$, extending $\Lambda$ and thus conflicting with $\Lambda^{\prime}$, $V^{3\Delta,j}_{\Lambda^{\prime}}$ does not include any log from $v_{k}$. Therefore, the senders of $V^{\Delta,j}_{\Lambda^{\prime}}\cap V^{3\Delta,j}_{\Lambda^{\prime}}$ and $V^{\Delta,i}_{\Lambda}\cap V^{3\Delta,i}_{\Lambda}$ are disjoint. Moreover, the senders of logs in $V^{\Delta,j}_{\Lambda^{\prime}}\cap V^{3\Delta,j}_{\Lambda^{\prime}}$ are by definition contained in $S^{3\Delta,j}$. Finally, the senders of logs in $V^{\Delta,i}_{\Lambda}$ are also all contained in $S^{3\Delta,j}$, because $v_{i}$ forwards them at time $\Delta$ and, by time $2\Delta$,  $v_{j}$ accepts either them or equivocation evidence for their senders.
@@
For the Graded Delivery property we show that an honest validator $v_{i}$ outputting $(\Lambda,1)$ implies that any honest validator $v_{j}$ participating in the output phase for grade 00 outputs $(\Lambda,0)$. At time $\Delta$, validator $v_{i}$ forwards all messages in $V^{\Delta,i}$. At time $2\Delta$, validator $v_{j}$ has equivocation evidence for the sender of any log in $V^{\Delta,i}\setminus V^{2\Delta,j}$, since otherwise such a log would also be contained in $V^{2\Delta,j}$. This equivocation evidence is forwarded by $v_{j}$ and received by $v_{i}$ by time $3\Delta$, so the senders of logs in $V^{\Delta,i}\setminus V^{2\Delta,j}$ are all considered as equivocators by $v_{i}$ then. This implies that $V^{\Delta,i}\setminus V^{2\Delta,j}$ and $V^{3\Delta,i}$ are disjoint, and thus so are $V^{\Delta,i}\cap V^{3\Delta,i}$ and $V^{\Delta,i}\setminus V^{2\Delta,j}$. Therefore, $V^{\Delta,i}\cap V^{3\Delta,i}\subseteq V^{2\Delta,j}$. Moreover, $S^{2\Delta,j}\subseteq S^{3\Delta,i}$, given that $v_{i}$ forwards at least a $\mathsf{LOG}$ message for each sender in $S^{2\Delta,j}$ by time $2\Delta$, and by time $3\Delta$ validator $v_{i}$ receives either the forwarded $\mathsf{LOG}$ messages or equivocation evidence for each sender. Validator $v_{i}$ outputs $(\Lambda,1)$ if $|V^{\Delta,i}_{{\Lambda}}\cap V^{3\Delta,i}_{{\Lambda}}|>|S^{3\Delta,i}|/2$, in which case we also have $|V^{2\Delta,j}_{{\Lambda}}|\geq|V^{\Delta,i}_{{\Lambda}}\cap V^{3\Delta,i}_{{\Lambda}}|>|S^{3\Delta,i}|/2\geq|S^{2\Delta,j}|/2$. Thus, if validator $v_{i}$ outputs $(\Lambda,1)$, then validator $v_{j}$ outputs $(\Lambda,0)$.
@@
For the Validity property, consider an honest validator $v_{i}\in H_{\Delta}\cap H_{3\Delta}$ participating in the output phase for grade 1, i.e., $v_{i}$ awake at $t=\Delta$ and $t=3\Delta$. Suppose that all validators $H_{0}\setminus{B_{3\Delta}}$ (initially awake and honest throughout the protocol) send logs that extends $\Lambda$. Since such validators are honest throughout the protocol, they never equivocate, so $V^{\Delta,i}_{\Lambda}$ and $V^{3\Delta,i}_{\Lambda}$ contain all of these logs. Therefore, $|V^{\Delta,i}_{\Lambda}\cap V^{3\Delta,i}_{\Lambda}|\geq|H_{0}\setminus{B_{3\Delta}}|=|H_{0,0,3\Delta}|$. $H_{0}\cup B_{3\Delta}$ contains all validators which might ever send a log during the whole protocol, so $|S^{3\Delta,i}|\leq|H_{0}\cup B_{3\Delta}|=|H_{0}\setminus{B_{3\Delta}}|+|B_{3\Delta}|=|H_{0,0,3\Delta}|+|B_{3\Delta}|$. Since by assumption $|H_{0,0,3\Delta}|>|B_{3\Delta}|$, we have that $|H_{0,0,3\Delta}|>|S^{3\Delta,i}|/2$, and thus $|V^{\Delta,i}_{\Lambda}\cap V^{3\Delta,i}_{\Lambda}|\geq|H_{0,0,3\Delta}|>|S^{3\Delta,i}|/2$. It follows that $v_{i}$ outputs $(\Lambda,1)$. The argument for grade 00 is nearly identical, and we omit it.
@@
Integrity follows from the fact that outputting $\Lambda$ with any grade requires a majority of unique log senders to have sent a log that extends $\Lambda$ and to not have equivocated. Moreover, we have $|H_{0}|\geq|H_{0,0,3\Delta}|>|B_{3\Delta}|$, and the condition $|H_{0}|>|B_{3\Delta}|$ ensures that any majority must contain at least a validator from $H_{0}$. Consider an honest validator $v_{i}$ which outputs $\Lambda$ with grade 00, after observing $|V^{2\Delta,i}_{\Lambda}|>|S^{2\Delta,i}|/2$ at time $2\Delta$, though no validator in $H_{0}$ has sent a log $\Lambda^{\prime}\succeq\Lambda$ at time $t=0$. So, $V^{2\Delta,i}_{\Lambda}$ does not contain any log received from validators in $H_{0}$. However, both the validators in $V^{2\Delta,i}_{\Lambda}$ and the ones in $H_{0}$ are counted as senders in $S^{2\Delta,i}$, and therefore, we have $|S^{2\Delta,i}|\geq|V^{2\Delta,i}_{\Lambda}|+|H_{0}|$. Note that $2|H_{0}|>|H_{0}|+|B_{3\Delta}|\geq|S^{2\Delta,i}|$, so $|H_{0}|>|S^{2\Delta,i}|/2$. Then, $|V^{2\Delta,i}_{\Lambda}|>|S^{2\Delta,i}|/2$ and $|S^{2\Delta,i}|\geq|V^{2\Delta,i}_{\Lambda}|+|H_{0}|$ together imply $|S^{2\Delta,i}|>|S^{2\Delta,i}|/2+|H_{0}|>|S^{2\Delta,i}|/2+|S^{2\Delta,i}|/2$, i.e., $|S^{2\Delta,i}|>|S^{2\Delta,i}|$, a contradiction. The argument for grade $1$ is almost identical, and we omit it.

### v3

@@ 5 TOB-SVD
The protocol still relies on the key ideas of the time-shifted quorum technique presented above. A notable difference is that grade 0 outputs in MR [12] are computed at time $3\Delta$ using $\mathsf{VOTE}$ messages, whereas our GA protocol does not have any other message other than $\mathsf{LOG}$ messages, and it computes grade 0 outputs at time $2\Delta$. That said, votes in their GA are cast (almost) in the same way as grade 0 outputs are computed in our protocol, and they are essentially used to propagate the information forward until the end of the protocol. Another difference is the use of $|V_{\Lambda}^{\Delta}\cap V_{\Lambda}^{{3\Delta}}|$ when determining outputs of grade $1$, which is related to the treatment of equivocations. Let us initially ignore that, and pretend that we output $(\Lambda,1)$ when $|V_{\Lambda}^{\Delta}|>|S^{3\Delta}|/2$.
@@
To ensure that the supporting $\mathsf{LOG}$ messages used when attempting to output a log $\Lambda$ with grade 0 are more than when doing so with grade $1$, we use $V^{\Delta}_{\Lambda}\cap V^{3\Delta}_{\Lambda}$ instead of just $V^{\Delta}_{\Lambda}$. Crucially, any validator $v_{k}$ which is seen as an equivocator by validator $v_{j}$ at time $2\Delta$ will also be seen as an equivocator by validator $v_{i}$ at time $3\Delta$, since $v_{j}$ would forward the equivocating $\mathsf{LOG}$ messages of $v_{k}$ upon receiving them. In particular, logs sent by $v_{k}$ will not be contained in $V^{3\Delta,i}$, and thus also not in $V^{\Delta,i}_{\Lambda}\cap V^{3\Delta,i}_{\Lambda}$. This ensures that $V^{\Delta,i}_{\Lambda}\cap V^{3\Delta,i}_{\Lambda}\subseteq V^{2\Delta,j}_{\Lambda}$, since any log in $V^{\Delta,i}\setminus V^{2\Delta,j}$ is an equivocation, which implies that it is absent from $V^{3\Delta,i}$. In other words, we separate the initial determination of possible supporting $\mathsf{LOG}$ messages from the final determination of which $\mathsf{LOG}$ messages should be treated as equivocations and removed, and we apply the time-shifted quorum technique to the set of equivocators as well: like the perceived participation level, the set of equivocators increases when going from the output phase for grade 0 to the output phase for grade $1$.
5.2 Graded Agreement with $k=3$ grades.
@@
Building upon our Graded Agreement protocol with $k=2$ grades, we extend it to a Graded Agreement with $k=3$ grades by applying the time-shifted quorum technique twice. The first application happens during time $2\Delta,4\Delta]$, and ensures the Graded Delivery property between grades 0 and $1$, exactly in the same way as in our Graded Agreement with $k=2$ grades. This application is nested inside the second one, which ensures the Graded Delivery property between grades $1$ and $2$, and happens during time $\Delta,5\Delta]$. Overall, the relevant inclusions are $V^{\Delta}_{\Lambda}\cap V^{5\Delta}_{\Lambda}\subseteq V^{2\Delta}_{\Lambda}\cap V^{4\Delta}_{\Lambda}\subseteq V^{3\Delta}_{\Lambda}$, for any $\Lambda$, and $S^{3\Delta}\subseteq S^{4\Delta}\subseteq S^{5\Delta}$ (each set can belong to a different validator). The protocol is given in Figure 2. This protocol lasts $5\Delta$ time instead of $3\Delta$ and, similarly to the previous GA, it requires that the number of Byzantine validators is less than half the number of active validators. Therefore, it in particular works in the $(5\Delta,0,\frac{1}{2})$-sleepy model.
@@
Upon input $\Lambda$, a validator $v_{i}$ runs the following algorithm whenever awake. All validators awake at time $3\Delta$ participate in the output phase for grade 0. A validator participates in the output phase for grade $1$ at time $4\Delta$ if and only if it was awake also at time $2\Delta$. A validator participates in the output phase for grade $2$ at time $5\Delta$ if and only if it was awake also at time $\Delta$. At any time, honest validators forward any message received. Up to two different $\mathsf{LOG}$ messages per sender are forwarded upon reception. 1. Input phase, $(t=0)$: Broadcast ${\langle\mathsf{LOG},\Lambda\rangle}_{i}$. 2. $(t=\Delta)$: Store $V^{\Delta}$. 3. $(t=2\Delta)$: Store $V^{2\Delta}$. 4. Output phase for grade 0, $(t=3\Delta)$: If $|V^{3\Delta}_{\Lambda}|>|S^{3\Delta}|/2$, output $(\Lambda,0)$. 5. Output phase for grade $1$, $(t=4\Delta)$: If awake at time $2\Delta$: If $|V^{2\Delta}_{\Lambda}\cap V^{4\Delta}_{\Lambda}|>|S^{4\Delta}|/2$, output $(\Lambda,1)$. 6. Output phase for grade $2$, $(t=5\Delta)$: If awake at time $\Delta$: If $|V^{\Delta}_{\Lambda}\cap V^{5\Delta}_{\Lambda}|>|S^{5\Delta}|/2$, output $(\Lambda,2)$.
@@
5.3 TOB-SVD
We conclude this section with our main result, TOB-SVD, a dynamically available Total-Order Broadcast protocol. Byzantine Total-Order Broadcast protocols can be built using Graded Agreement primitives, with some protocols invoking Graded Agreement multiple times for a single decision [12, 10, 11]. At its core, dynamically available TOB aims to guarantee the total order delivery of messages among honest validators, even in the face of challenges like dynamic participation. However, existing TOB protocols [12, 10, 11, 9] have grappled with challenges related to latency [12], resilience [10], and scalability [9]. Multiple invocations of GA for a single decision, meaning multiple voting phases per new block, might exacerbate scalability and latency issues, especially when considering implementation in blockchain networks with hundreds of thousands of validators weighing in on every decision.
@@ 5.3 TOB-SVD
Our protocol, which works in the $(5\Delta,2\Delta,\frac{1}{2})$-sleepy model, proceeds over a series of views, each spanning a duration of $4\Delta$. Every view $v$ initiates a Graded Agreement $GA_{v}$ with grades 0, 11, and $2$, which extends and overlaps with the following $GA_{v+1}$ during view $v+1$. This structure implies that a single view does not encapsulate a full cycle of a Graded Agreement; instead, a $GA_{v}$ initiated in view $v$ concludes its operations only in the succeeding view $v+1$.
@@
Figure 3: In the middle, views v1v-1, $v,$ and $v+1$ of our Total-Order Broadcast protocol, each with its three phases. At the top and bottom, respectively, $GA_{v-1}$ and $GA_{v}$ — the Graded Agreement instances that are run as part of the TOB protocol. Arrows indicate that outputs of a GA are used by a parallel phase of the TOB and/or the next GA: outputs of grade 0 of $GA_{v-1}$ are extended by proposals of view $v$, outputs of grade $1$ of $GA_{v-1}$ are extended by votes in the TOB, which exactly correspond to the inputs to $GA_{v}$, while outputs of grade $2$ of $GA_{v-1}$ are decided in view $v$.
Specifically, the protocol, which is presented in Figure 4, proceeds in views of $4\Delta$ time each. We let $t_{v}=4\Delta v$ be the beginning of view $v$. To each view $v$ corresponds a Graded Agreement $GA_{v}$, which runs in the time interval $t_{v}+\Delta,t_{v}+6\Delta]=[t_{v}+\Delta,t_{v+1}+2\Delta]$, i.e., $GA_{v}$ takes up some of view $v+1$ as well. Moreover, the GA invocations are not perfectly sequential, as $GA_{v}$ and $GA_{v+1}$ overlap during time $t_{v+1}+\Delta,t_{v+1}+2\Delta]$ (Figure 3).
At the beginning of each view $v$ there is a proposal phase corresponding to the output phase for grade 0 of $GA_{v-1}$. The leader of view $v$ proposes a log $\Lambda$ extending its grade 0 output (a candidate), if it has one.
This is followed by a decision phase at time $t_{v}+2\Delta$, where logs proposed (at best) in the previous view (v1v-1) can be decided. In particular, this phase corresponds to the output phase for grade $2$ of $GA_{v-1}$, and such grade $2$ outputs are decided. This is safe, because Graded Delivery of Graded Agreement ensures that all honest validators (among those participating in the output phase for grade $1$) output with grade $1$ a decided log, thus locking on it and therefore inputting to $GA_{v}$ a log extending it. During the decision phase, validators also update $V$ for $GA_{v}$ with all $\mathsf{LOG}$ messages received until that point in time for $GA_{v}$ from non-equivocating validators, as part of the ongoing $GA_{v}$.
Finally, at time $t_{v}+3\Delta$, only the action required by $GA_{v}$ is performed, i.e., to store in $V^{2\Delta}$ all $\mathsf{LOG}$ messages received until that point from non-equivocating validators.
Outputs of $GA_{-1}$, which are used in view $v=0$, are all formally defined to be the log containing only the Genesis log. In each view $v\geq 0$, awake validators participate in the GA instances that are ongoing, and in addition behave as specified here whenever they have the required GA outputs to do so. Validators do not perform actions which require outputs they do not have. 1. Propose ($t=t_{v}$): Output phase for grade 0 of $GA_{v-1}$. Propose $\Lambda^{\prime}$ extending $\Lambda$, the highest log output with grade 0 by $GA_{v-1}$ (candidate), accompanied by the VRF value for view $v$. 2. Vote ($t=t_{v}+\Delta$): Output phase for grade $1$ of $GA_{v-1}$. $GA_{v}$ starts. Let $L_{v-1}$ be the highest log output with grade $1$ by $GA_{v-1}$ (lock). After discarding equivocating proposals, input to $GA_{v}$ the proposal with the highest VRF value extending $L_{v-1}$, or $L_{v-1}$ if no such proposal exists. 3. Decide ($t=t_{v}+2\Delta$): Output phase for grade $2$ of $GA_{v-1}$, which ends. Decide the highest log $\Lambda$ output with grade $2$ by $GA_{v-1}$. 4. ($t=t_{v}+3\Delta$): Do nothing (other than what is required by the ongoing $GA_{v}$).
@@
On the number of grades. Our protocol requires $k=3$ grades to reach a decision with a single instance of the GA. Specifically, grade 2 enables an instant decision, which would otherwise necessitate two instances of GA, each with $k=2$ grades. For comparison, in constructions like 1/3MMR, validators first input their logs into the first GA. Afterward, they input the longest log output with grade 1 from the first GA into the second GA. The decision is then made from the grade 1 output of the second GA, which corresponds to the grade 2 output in our GA with $k=3$ grades.
@@
It is important to note that the decision phase of our TOB protocol is solely focused on making decisions and nothing else. By omitting the decision phase, the protocol can be simplified, still maintaining one GA per decision. However, this modification would shift the protocol from ensuring deterministic Safety to probabilistic Safety. As a result, it would align more closely with a different line of research focused on probabilistically safe, dynamically available protocols, such as those presented in [5, 6]. However, exploring probabilistically safe TOB protocols falls outside the scope of this work.
6 Analysis
In this section, we analyze the protocols in Section 5.
@@ 6.1 Graded Agreement with $k=2$ grades
We begin by analyzing the Graded Agreement protocol with $k=2$ grades. Recall that this protocol works in the $(3\Delta,0,\frac{1}{2})$-sleepy model and that honest validators which are awake during the output phase for grade 1, at time $t=3\Delta$, only participate in it if they were also awake at time $t=\Delta$.
@@
Let $m=(\Lambda^{\prime\prime},v_{k})$ for some $\Lambda^{\prime\prime}\succeq\Lambda$ and some validator $v_{k}$, and say $m\in V^{\Delta,i}\cap V^{3\Delta,i}$. Since $m\in V^{\Delta,i}$, validator $v_{i}$ forwards message ${\langle\mathsf{LOG},\Lambda^{\prime\prime}\rangle}_{k}$ at time $\Delta$ and $v_{j}$ receives it by time $2\Delta$. Therefore, either $m\in V^{3\Delta,j}$ or $v_{k}\notin V^{3\Delta,j}$, because receiving any log other than $\Lambda^{\prime\prime}$ from $v_{k}$ would lead to $v_{j}$ having equivocation evidence for $v_{k}$. Since $m$ is for $\Lambda^{\prime\prime}$, extending $\Lambda$ and thus conflicting with $\Lambda^{\prime}$, $V^{3\Delta,j}_{\Lambda^{\prime}}$ does not include any log from $v_{k}$. Hence, the senders of $V^{\Delta,j}_{\Lambda^{\prime}}\cap V^{3\Delta,j}_{\Lambda^{\prime}}$ and $V^{\Delta,i}_{\Lambda}\cap V^{3\Delta,i}_{\Lambda}$ are disjoint. Moreover, the senders of logs in $V^{\Delta,j}_{\Lambda^{\prime}}\cap V^{3\Delta,j}_{\Lambda^{\prime}}$ are by definition contained in $S^{3\Delta,j}$. Finally, the senders of logs in $V^{\Delta,i}_{\Lambda}$ are also all contained in $S^{3\Delta,j}$, because $v_{i}$ forwards them at time $\Delta$ and, by time $2\Delta$,  $v_{j}$ accepts either them or equivocation evidence for their senders.
@@
For the Graded Delivery property we show that an honest validator $v_{i}$ outputting $(\Lambda,1)$ implies that any honest validator $v_{j}$ participating in the output phase for grade 0 outputs $(\Lambda,0)$. At time $\Delta$, validator $v_{i}$ forwards all messages in $V^{\Delta,i}$. At time $2\Delta$, validator $v_{j}$ has equivocation evidence for the sender of any log in $V^{\Delta,i}\setminus V^{2\Delta,j}$, since otherwise such a log would also be contained in $V^{2\Delta,j}$. This equivocation evidence is forwarded by $v_{j}$ and received by $v_{i}$ by time $3\Delta$, so the senders of logs in $V^{\Delta,i}\setminus V^{2\Delta,j}$ are all considered as equivocators by $v_{i}$ then. This implies that $V^{\Delta,i}\setminus V^{2\Delta,j}$ and $V^{3\Delta,i}$ are disjoint, and thus so are $V^{\Delta,i}\cap V^{3\Delta,i}$ and $V^{\Delta,i}\setminus V^{2\Delta,j}$. Therefore, $V^{\Delta,i}\cap V^{3\Delta,i}\subseteq V^{2\Delta,j}$. Moreover, $S^{2\Delta,j}\subseteq S^{3\Delta,i}$, given that $v_{i}$ forwards at least a $\mathsf{LOG}$ message for each sender in $S^{2\Delta,j}$ by time $2\Delta$, and by time $3\Delta$ validator $v_{i}$ receives either the forwarded $\mathsf{LOG}$ messages or equivocation evidence for each sender. Validator $v_{i}$ outputs $(\Lambda,1)$ if $|V^{\Delta,i}_{{\Lambda}}\cap V^{3\Delta,i}_{{\Lambda}}|>|S^{3\Delta,i}|/2$, in which case we also have $|V^{2\Delta,j}_{{\Lambda}}|\geq|V^{\Delta,i}_{{\Lambda}}\cap V^{3\Delta,i}_{{\Lambda}}|>|S^{3\Delta,i}|/2\geq|S^{2\Delta,j}|/2$. Thus, if validator $v_{i}$ outputs $(\Lambda,1)$, validator $v_{j}$ outputs $(\Lambda,0)$.
@@
For the Validity property, consider an honest validator $v_{i}\in H_{\Delta}\cap H_{3\Delta}$ participating in the output phase for grade 1, i.e., $v_{i}$ awake at $t=\Delta$ and $t=3\Delta$. Suppose that all the validators $H_{0}\setminus{B_{3\Delta}}$ (initially awake and honest throughout the protocol) send logs that extends $\Lambda$. Since such validators are honest throughout the protocol, they never equivocate, so $V^{\Delta,i}_{\Lambda}$ and $V^{3\Delta,i}_{\Lambda}$ contain all of these logs. Therefore, $|V^{\Delta,i}_{\Lambda}\cap V^{3\Delta,i}_{\Lambda}|\geq|H_{0}\setminus{B_{3\Delta}}|=|H_{0}\cup{B_{3\Delta}}|-|{B_{3\Delta}}|$. Note that $H_{0}\cup B_{3\Delta}$ contains all validators which might ever send a log during the whole protocol. Then, due to Condition (1) and $\rho=\frac{1}{2}$, $|V^{\Delta,i}_{\Lambda}\cap V^{3\Delta,i}_{\Lambda}|>|S^{3\Delta,i}|/2$. It follows that $v_{i}$ outputs $(\Lambda,1)$. The argument for grade 0 is nearly identical, and we omit it.
Integrity follows from $\rho=\frac{1}{2}$ and the fact that outputting $\Lambda$ with any grade requires a majority of unique log senders to have sent a log that extends $\Lambda$ and to not have equivocated.
@@
Consider an honest validator $v_{i}$ which outputs $\Lambda$ with grade 0, after observing $|V^{2\Delta,i}_{\Lambda}|>|S^{2\Delta,i}|/2$ at time $2\Delta$, though no validator in $H_{0}$ has sent a log $\Lambda^{\prime}\succeq\Lambda$ at time $t=0$. So, $V^{2\Delta,i}_{\Lambda}$ does not contain any log received from validators in $H_{0}$. However, both the validators in $V^{2\Delta,i}_{\Lambda}$ and the ones in $H_{0}$ are counted as senders in $S^{2\Delta,i}$, and therefore, we have $|S^{2\Delta,i}|\geq|V^{2\Delta,i}_{\Lambda}|+|H_{0}|$. Note that $|H_{0}\cup B_{3\Delta}|\geq|S^{2\Delta,i}|$. Then, due to Condition (1) and $\rho=\frac{1}{2}$, $|H_{0}|\geq|H_{0}\cup B_{3\Delta}|-|B_{3\Delta}|>|H_{0}\cup B_{3\Delta}|/2\geq|S^{2\Delta,i}|/2$. Then, $|V^{2\Delta,i}_{\Lambda}|>|S^{2\Delta,i}|/2$ and $|S^{2\Delta,i}|\geq|V^{2\Delta,i}_{\Lambda}|+|H_{0}|$ together imply $|S^{2\Delta,i}|>|S^{2\Delta,i}|/2+|H_{0}|>|S^{2\Delta,i}|/2+|S^{2\Delta,i}|/2$, i.e., $|S^{2\Delta,i}|>|S^{2\Delta,i}|$, a contradiction. The argument for grade $1$ is almost identical, and we omit it.